# Estimating Unknown Population Sizes Using the Hypergeometric Distribution

### v1

@@ 4.2 Benchmarks
Having demonstrated that the inference problem is tractable, we turn to our complete latent-variable model. To show the advantages of using the hypergeometric distribution for inference, we compare it to other popular distributional assumptions.
Multinomial: The multinomial likelihood is commonly used in recommender systems - while it cannot be used to estimate counts, it can used to perform unsupervised clustering. Discrete count data must be normalized to produce a distribution over features (Liang et al., 2018).
@@ 4.2.1 Count Estimation Experiment
Since we are proposing a generative model for count data, we first show that our VAE with hypergeometric likelihood is able to generate a better estimate of the data than competing likelihoods. We simulate datasets with K = 100, 1000, or 10,000 categories, each consisting of a mixture of either 3 or 10 discrete distributions (simulation details in Appendix A). We train each the base model with each likelihood until convergence and then quantify the estimation error as the discrepancy between the ground truth distribution and the model’s estimate, using both mean absolute error (MAE) and median percentage error (MPE) metrics.
@@ 4.2.2 Clustering Experiment
We argue that the hypergeometric likelihood is able to correctly infer the latent structure in cases where other likelihood functions fail. A major benefit of using a latent variable model $p({\bm{c}},{\bm{z}})$ is that we can make use of the learned latent representation ${\bm{z}}$ to perform deep unsupervised clustering. Both the multinomial and Poisson likelihoods have been used to learn useful latent representations of count data. The assumptions of the hypergeometric distribution, which considers both total counts and count fraction, can be thought of as a superset of the assumptions of the multinomial distribution, which only considers count fraction. For example, this difference would become important in datasets composed of a mixture of distributions with similar count fractions but different total counts.
@@
Representative results for this experiment are again shown in Table 1. Our method outperforms all other likelihoods in its ability to learn a latent space that is able to differentiate between distributions. In particular, it is the only examined method able to distinguish between count distributions with similar probability distributions over features. These results are consistent across a wide variety of simulated datasets (Appendix D).
@@ 4.3 Application to Reading Passage Complexity
In the field of NLP, it is valuable to be able to measure the linguistic complexity of a text. For example, quantifying the readability of text passages is essential in the education context, as it allows texts to be correctly matched to students at the appropriate level of reading skill. We hypothesize that any short text passage can be thought of as a bag-of-words (BoW) that is sampled from a larger latent BoW, and that the size and complexity of this underlying BoW provides a better measure of the readability of a passage than the passage itself. That is, a passage that is generated from a richer underlying vocabulary will demand a higher level of reading comprehension from its reader.
We test this hypothesis using the CommonLit Ease of Readability (CLEAR) Corpus (Crossley et al., 2023), an open-source dataset consisting of almost 5000 text excerpts sourced from Grade 3-12 reading curricula. Numerous metrics have been proposed to estimate the complexity of reading passages, and each excerpt has been annotated with a number of well-established readability indices, including those can be automatically calculated by ARTE, the Automatic Readability Tools for English (Choi & Crossley, 2022). Additionally, each passage has a Bradley-Terry reading ease score derived from the qualitative assessment by teachers of the relative reading difficulty between pairs of passages.
This scenario presents an ideal application of our proposed hypergeometric method, which can capture both the dependence between the choice of words (the selection of one word comes at the cost of its synonyms, and at the benefit of words from the same topic) and the vocabulary size as measured in token counts. Additionally, since there is no clear classification of the texts, a continuous latent variable model is well suited. We create a bag-of-words representation by tokenizing the 4724 text excerpts in the CLEAR Corpus, producing a count matrix over the resulting tokens (Appendix C for details). We then train our unsupervised model on this BoW representation of the excerpts, which enables us to infer a latent BoW that underlies each passage. We use this estimate as a proxy for passage readability using both the total estimated number of tokens $t_{total}$ and the number of unique tokens $t_{unique}$ (tokens with non-zero count) in each passage’s latent BoW. We then calculate the Pearson correlation across the 4724 passages between each of our proposed metrics ($t_{total}$ and $t_{unique}$) and the readability indices provided in the CLEAR dataset, in order to assess the association strength.
@@
Similarly, Figure 3(b) shows that the vocabulary size of the latent BoW (the number of unique non-zero tokens for an excerpt) inferred by our model also aligns with the complexity of the passage much better than if we used the original text’s vocabulary. These results demonstrate that our method is able to infer a latent vocabulary underlying each reading passage that convincingly reflects the independently-assessed readability of the text.
@@ 4.4 Application to Single-Cell Genomics
The detailed measurement of the contents of individual cells promises to vastly improve our understanding of fundamental biology. In recent years, the advancement of high-throughput techniques in the field of single-cell genomics has enabled the collection of large numbers of gene transcripts from individual cells, resulting in vast count matrices. Each cell has a finite population of transcripts that can be captured - on the order of one hundred thousand to a million - and a transcript can only be captured once prior to sequencing. It is accepted that the main source of technical noise in this data is due to under-sampling (Kuo et al., 2023), leading to the well-known phenomenon of dropout (inflated occurrence of zeros in the final count matrix). This technical noise hinders scientists’ ability to draw meaningful scientific conclusions from these experiments.
@@
Because there is typically no way of knowing the true number of transcripts of each gene in a given cell, we focus on what is known as a spike-in experiment, which does provide a ground-truth. In the experiment we consider (Ziegenhain et al., 2022), a known concentration of a solution of synthetic RNA is placed in small wells, and human cells are individually placed in a subset of these wells. After the transcripts present in the wells are captured and sequenced, we obtain experimental counts corresponding to both the human RNA (for which the true amount is unknown) and the synthetic RNA (true amount is known). The wells that did not have cells in them should have an equal amount of the synthetic RNA across measurements, so we can therefore use the measured counts of synthetic RNA in both empty and cell-containing wells as a ground-truth reference to evaluate our estimated human gene transcript counts.
@@
Figure 5 compares the distribution of measured counts in this dataset for one specific synthetic RNA (#12) in the empty and HEK293T measurements. This exemplifies the stochastic under-sampling that occurs, as the true counts should be identical across all observations - approximately 4000 for this particular spike-in RNA. It is also clear that the presence of human RNA in the HEK293T observations significantly lowers the amount of synthetic RNA that is captured, which supports the assumption that there is dependence between counts, i.e. a transcript is captured at the cost of another.
@@
We train our hypergeometric model on the SPIKE dataset to infer the ground-truth count matrix. We use the top 10,000 genes (categories) by mean transcript count, for computational efficiency and because many genes are not expressed in this celltype. Model and training hyperparameters are given in Appendix B. Summary results of the final estimated count matrix are shown in Figure 6.
@@
Figure 5(a) shows the inferred total number of counts per observation (the sum of the estimated count matrix rows). The authors estimate the true total amount of synthetic RNA per observation as approximately 30,000, which closely aligns with our model estimate. The unknown total amount of RNA in the human kidney cells is estimated by our model to have a median of 260,000. Note that the number of observations available in this dataset is very small compared to typical single-cell datasets (without ground-truth), and this estimate would likely be even more accurate and have a lower variance with a larger number of observed cells. Figure 5(b) shows the estimated counts specifically for spike-in #12, whose ground-truth value is approximately 4000. We can see that although the initial distribution of counts for this synthetic RNA has a significantly different mean in the empty and cell-containing observations, our model produces an estimate near the true value, and brings both distributions into alignment (this is also true in the other spike-in RNA). These results show that we are able to recover the known ground-truth distribution of synthetic RNA that has been corrupted by the stochastic capture process.

### v2

@@ 4.2 Benchmarks
Having demonstrated that the inference problem is tractable, we turn to our complete latent-variable model. To show the advantages of using the hypergeometric distribution for inference, we compare it to other common distributional assumptions.
Multinomial: The multinomial likelihood is commonly used in recommender systems - while it cannot be used to infer absolute counts, it can used to perform unsupervised clustering and estimate relative category frequency (Liang et al., 2018).
@@ 4.2.1 Count Estimation Experiment
Since we are proposing a generative model for count data, we first show that our VAE with hypergeometric likelihood is able to better estimate the data than using alternate likelihoods. We simulate datasets with K = 100, 1000, or 10,000 categories, each consisting of a mixture of either 3 or 10 discrete distributions (simulation details in Appendix A). We train the base model with each likelihood until convergence and then quantify the estimation error as the discrepancy between the ground truth distribution and the model’s estimate, using both mean absolute error (MAE) and median percentage error (MPE) metrics.
@@ 4.2.2 Clustering Experiment
We next show that the hypergeometric likelihood can correctly infer the latent structure in cases where other likelihood functions fail. A major benefit of using a latent variable model is that we can make use of the learned latent representation ${\bm{z}}$ to perform deep unsupervised clustering (Lim et al., 2020). Both the multinomial and Poisson likelihoods have been used to learn useful latent representations of count data (Bouguila, 2010). The assumptions of the hypergeometric distribution, which considers both absolute and relative counts, can be thought of as a superset of the assumptions of the multinomial distribution, which only considers relative counts. For example, this difference would become important in datasets composed of a mixture of distributions with similar count fractions but different total counts.
@@
Representative results for this experiment are again shown in Table 3. Our method outperforms all other likelihoods in its ability to learn a latent space that is able to differentiate between underlying distributions. In particular, it is the only examined method able to distinguish between count distributions with similar probability distributions over features. These results are consistent across a wide variety of simulated datasets (Appendix D).
@@ 4.3 Application to Reading Passage Complexity
In the field of NLP, it is valuable to be able to measure the linguistic complexity of a text. For example, quantifying the readability of text passages is essential in the education context, as it allows texts to be correctly matched to students at the appropriate level of reading skill (Sarti et al., 2021). We hypothesize that any short text passage can be thought of as a bag-of-words (BoW) that is sampled from a larger latent BoW, and that the size and complexity of this underlying BoW provides a better measure of the readability of a passage than the passage itself. That is, a text passage that is generated from a richer underlying vocabulary will demand a higher level of reading comprehension from its reader.
We test this hypothesis using the CommonLit Ease of Readability (CLEAR) Corpus (Crossley et al., 2023), an open-source dataset consisting of almost 5000 text excerpts sourced from Grade 3-12 reading curricula. Numerous metrics have been proposed to estimate the complexity of reading passages, and each excerpt has been annotated with a number of well-established readability indices, including those can be automatically calculated by ARTE, the Automatic Readability Tools for English (Choi & Crossley, 2022). Additionally, each passage has a Bradley-Terry reading ease score (Bradley & Terry, 1952) derived from the qualitative assessment by teachers of the relative reading difficulty between pairs of passages.
This scenario presents an ideal application of our proposed hypergeometric method, which can capture both the dependence between the choice of words (the selection of one word comes at the cost of its synonyms, and to the benefit of words from the same topic) and the vocabulary size as measured in token counts. Additionally, since there is no clear classification of the texts, a continuous latent variable model is well suited. We create a bag-of-words representation by tokenizing the 4724 text excerpts in the CLEAR Corpus, producing a count matrix over the resulting tokens (Appendix C for details). We then train our unsupervised model on this BoW representation of the excerpts, which enables us to infer a latent BoW that underlies each passage. We use this estimate as a proxy for passage readability using both the total estimated number of tokens $t_{total}$ and the number of unique tokens $t_{unique}$ (tokens with non-zero count) in each passage’s latent BoW. We then calculate the Pearson correlation across the 4724 passages between each of our proposed metrics ($t_{total}$ and $t_{unique}$) and the readability indices provided in the CLEAR dataset, in order to assess the association strength.
@@
Similarly, Figure 4(b) shows that the vocabulary size of the latent BoW (the number of unique non-zero tokens for a passage) inferred by our model also aligns with the complexity of the passage much better than if we used the original text’s vocabulary. These results demonstrate that our method is able to infer a latent vocabulary underlying each reading passage that convincingly reflects the independently-assessed readability of the text.
@@ 4.4 Application to Single-Cell Genomics
The detailed measurement of the contents of individual cells promises to vastly improve our understanding of fundamental biology. In recent years, the advancement of high-throughput techniques in the field of single-cell genomics has enabled the collection of large numbers of gene transcripts from individual cells, resulting in vast count matrices. Each cell has a finite population of transcripts that can be captured - on the order of one hundred thousand to a million - and a transcript can only be captured once prior to sequencing (De Simone et al., 2019). It is accepted that the main source of technical noise in this data is due to under-sampling (Kuo et al., 2023), leading to the well-known phenomenon of dropout (inflated occurrence of zeros in the final count matrix). This technical noise hinders scientists’ ability to draw meaningful scientific conclusions from these experiments.
@@
Because there is typically no way of knowing the true number of transcripts of each gene in a given cell, we focus on what is known as a spike-in experiment, which does provide a ground-truth. In the experiment we consider (Ziegenhain et al., 2022), a known concentration of a solution of synthetic RNA is placed in small wells, and human cells are individually placed in a subset of these wells. After the transcripts present in the wells have been captured and sequenced, we obtain experimental counts corresponding to both the human RNA (for which the true amount is unknown) and the synthetic RNA (true amount is known). The wells that did not have cells in them should have an equal amount of the synthetic RNA across measurements, so we can therefore use the measured counts of synthetic RNA in both empty and cell-containing wells as a ground-truth reference to evaluate our estimated human gene transcript counts.
@@
Figure 5 compares the distribution of measured counts in this dataset for one specific synthetic RNA (#12) for the empty and HEK293T measurements. This difference between these distributions exemplifies the stochastic under-sampling that occurs, as the true counts should be identical across all observations - approximately 4000 for this particular spike-in RNA. It is also clear that the presence of the additional human RNA in the HEK293T observations significantly lowers the amount of synthetic RNA that is captured, which supports the assumption that there is dependence between counts, i.e. a transcript is captured at the cost of another.
@@
We train our hypergeometric model on the SPIKE dataset to infer the ground-truth count distribution. We use the top $K=10,000$ genes (categories) by mean transcript count, for computational efficiency and because many genes are not expressed in this celltype. Model and training hyperparameters are given in Appendix B. Summary results of the final estimated count matrix are shown in Figure 6.
@@
Figure 6(a) shows the inferred total number of counts per measurement well (the sum of the estimated count matrix rows). We can see that the variance of the total number of spike-in RNA, which is known to be constant, is significantly reduced in our estimate. The experiment’s authors estimated the total amount of synthetic RNA per observation as approximately 30,000, which aligns with our model estimate. The unknown total amount of RNA in the human kidney cells is estimated by our model to have a median of 260,000, in line with a previous estimate of 200,000 (Shapiro et al., 2013). Note that the number of observations available in this dataset is very small compared to typical single-cell datasets (without ground-truth), and this estimate is expected to be even more accurate and have a lower variance with a larger number of observed cells. Figure 6(b) shows the estimated counts specifically for spike-in #12, whose ground-truth value is approximately 4000. We can see that although the initial distribution of counts for this synthetic RNA has a significantly different mean in the empty and cell-containing observations, our model produces an estimate near the true value, and brings both distributions into alignment (this is also true in the other spike-in RNA). These results show that we are able to recover the known ground-truth distribution of synthetic RNA that has been corrupted by the stochastic capture process.
@@ 5 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of machine learning, in particular through its applications to genomics experimental data.